### v3

# Automorphisms of free metabelian Lie algebras, I

Abstract.
We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame. We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions. The same results hold for fields of any characteristic when $n\geq 5$. These results contradict some long-standing results in the area.
We also prove that a large class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
Ualbai Umirbaev [^1] Department of Mathematics, Wayne State University, Detroit, MI 48202, USA; Department of Mathematics, Al-Farabi Kazakh National University, Almaty, 050040, Kazakhstan; and Institute of Mathematics and Mathematical Modeling, Almaty, 050010, Kazakhstan, e-mail: umirbaev@wayne.edu
Mathematics Subject Classification (2020): 17B40, 17B01, 17B30, 16W20.
Key words: Automorphism, derivation, free metabelian Lie algebra.
1. Introduction
The well known Jung–van der Kulk Theorem [16, 20] states that every automorphism of the polynomial algebra $K[x,y]$ in two variables $x,y$ over an arbitrary field $K$ is tame. An analogue of this result for free associative algebras in two variables was proven by Czerniakiewicz and Makar-Limanov [12, 24]. The same result is true for free Poisson algebras in two variables over a field of characteristic zero [26] (see also [25, 27, 65]). Moreover, the automorphism groups of polynomial algebras, free associative algebras, and free Poisson algebras in two variables are isomorphic. The tameness of automorphisms of free right-symmetric algebras in two variables was proven in [18].
For any free algebra $A$ in the free variables $x_{1},\ldots,x_{n}$ denote by $\phi=(f_{1},\ldots,f_{n})$ the endomorphism of $A$ determined by $\phi(x_{i})=f_{i}$ for all $1\leq i\leq n$.
The automorphism groups of commutative and associative algebras generated by three elements are much more complicated. The well-known Nagata automorphism [31]
$(x+2y(zx-y^{2})+z(zx-y^{2})^{2},y+z(zx-y^{2}),z)$
of the polynomial algebra $K[x,y,z]$ over a field $K$ of characteristic 0 is proven to be wild [46, 60]. The well-known Anick automorphism (see [11, p. 398])
$(x+z(xz-zy),\,y+(xz-zy)z,\,z)$
of the free associative algebra $K\langle x,y,z\rangle$ over a field $K$ of characteristic 0 is also proven to be wild [61, 63]. The Nagata automorphism gives an example of a wild automorphism of free Poisson algebras in three variables. Recently Shestakov and Zhang [48] constructed an analogue of the Anick automorphism for free Poisson algebras in three variables.
It is well known [55] that the Nagata and Anick automorphisms are stably tame, that is, they become tame after adding one more variable.
A classical theorem of combinatorial group theory asserts that every subgroup of a free group is free. Nielsen established this remarkable result in 1921 for finitely generated subgroups [37]; in 1927, it was extended to all subgroups by Schreier [52]. A variety of universal algebras is called Nielsen-Schreier, if any subalgebra of a free algebra of this variety is free. In 1953 Shirshiv and in 1956 Witt proved that [49, 67] the varieties of Lie algebras and Lie $p$-algebras over a field are Nielsen-Schreier. The varieties of all non-associative algebras [21], commutative and anti-commutative algebras [50], Lie superalgebras [28, 54], and Lie $p$-superalgebras [29] over a field are also Nielsen-Schreier. It was recently shown [13] that the varieties of pre-Lie (also known as right-symmetric) algebras and Lie-admissible algebras over a field of characteristic zero are Nielsen-Schreier. Some other examples of Nielsen-Schreier varieties can be found in [9, 30, 45, 57, 59].
In 1924, Nielsen also proved [38] that the automorphism group of a finitely generated free group is generated by Nielsen transformations, and represented this group as a finitely defined group. In 1964 P. Cohn proved [10] that all automorphisms of finitely generated free Lie algebras over a field are tame. Later this result was extended to free algebras of Nielsen-Schreier varieties [22]. Defining relations between the elementary automorphisms of free algebras of Nielsen-Schreier varieties are given in [64].
One of the classical and well-studied varieties of groups is the variety of metabelian groups. In 1965 Bachmuth [2] proved that all automorphisms of the free metabelian group of rank $2$ are tame. The first example of a non-tame automorphism of the free metabelian group of rank $3$ was constructed by Chein [8]. Bachmuth-Mochizuki proved [3] that the group of automorphisms of the free metabelian group of rank $3$ is not even finitely generated. Strengthening these results, Roman’kov proved [41] that the free metabelian group of rank $3$ contains non-tame primitive elements and gave a method of constructing of such elements. The well known Bachmuth-Mochizuki-Roman’kov Theorem [4, 40] states that every automorphism of the free metabelian group of rank $n\geq 4$ is tame. More information about automorphisms of free metabelian groups can be found in [44].
The variety of metabelian Lie algebras is also a classical and well-studied variety of Lie algebras. Throughout this paper, $M_{n}$ denotes the free metabelian Lie algebra of rank $n$ over a field $K$ generated by $x_{1},\ldots,x_{n}$. It is well known that any nontrivial exponential automorphism of $M_{2}$ is wild (see, for example [1, 51]). In 1992, Drensky [14] proved that the exponential automorphism $\mathrm{exp}(\mathrm{ad}[x_{1},x_{2}])$ of $M_{3}$ is wild. More examples of wild automorphisms of $M_{3}$ are given in [17, 32, 43].
In 1992 Bahturin and Nabiyev [5] claimed that every nontrivial exponential automorphism of the free metabelian Lie algebra $M_{n}$ of any rank $n\geq 2$ over a field of characteristic zero is wild. More examples of wild automorphisms of $M_{n}$ of rank $n\geq 4$ were given in 2008 by Özcurt and Ekici [39]. For a long time, the area specialists believed these results were true. Unfortunately, both articles have fatal errors [66].
We show that all Chein automorphisms (or one-row transformations) of lower degree $\geq 4$ of a free metabelian Lie algebra $M_{n}$ of rank $n\geq 4$ over an arbitrary field $K$ of characteristic $\neq 3$ are tame. We then show that all exponential automorphisms of $M_{n}$ of lower degree $\geq 5$ are also tame under the same conditions. The same results hold for fields of any characteristic when $n\geq 5$. These results contradict those of [5, 39].
We also prove that a large important class of automorphisms of $M_{n}$ of rank $n\geq 4$ that move only two variables are almost tame, that is, they can be expressed as a product of Chein automorphisms.
More precise formulations of the results are given in the following sections.
The paper is organized as follows. In Section 2 we define some important automorphisms of $M_{n}$ and give precise formulations some of our results. In Section 3, using a representation of $M_{n}$ via Magnus embedding and Fox derivatives, we describe Chein autoimorphisms of $M_{n}$. In Section 4 we prove that a huge part of Chein and exponential automorphisms are tame. In Section 5 we prove that a huge class of automorphisms that move only two variables are almost tame.
2. Automorphisms of $M_{n}$
Let $M_{n}$ be the free metabelian Lie algebra over a field $K$ with free generators $x_{1},\ldots,x_{n}$. The set of all right normed Lie monomials of the form
(1) $[\ldots[[x_{i},x_{i_{1}}],x_{i_{2}}],\ldots,x_{i_{k}}],\ i>i_{1}\leq i_{2}\leq\ldots\leq i_{k},$
is a linear basis of $M_{n}^{\prime}=[M_{n},M_{n}]$ (see, for example [6, p. 139]).
Let $\mathrm{Aut}(M_{n})$ be the group of all automorphisms of $M_{n}$. An automorphism of $M_{n}$ of the form
$(x_{1},x_{2},\ldots,x_{i-1},\alpha x_{i}+f,x_{i+1},\ldots x_{n}),$
where $0\neq\alpha\in K$ and $f$ does not contain $x_{i}$, is called elementary. The subgroup $\mathrm{TAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all elementary automorphisms is called the tame automorphism group, and the elements of this subgroup are called tame automorphisms of $M_{n}$. Nontame automorphisms of $M_{n}$ are called wild.
Defining relations among elementary automorphisms have been described for polynomial algebras in three variables [61, 62], for finitely generated free Lie algebras [64], and for free metabelian Lie algebras in three variables [33].
As usual, the group of all linear automorphisms of $M_{n}$ will be identified with the group of all invertible matrices $GL_{n}(K)$. For every pair of integers $s,t$ with $1\leq s\neq t\leq n$, let $(st)$ denote the automorphism that simply permutes the variables $x_{s}$ and $x_{t}$ and leaves all other variables unchanged. The group generated by all $(st)$ is isomorphic to the symmetric group $S_{n}$.
In 2015, Nauryzbaev [35] proved that the group of tame automorphisms of $M_{n}$ is generated by all linear automorphisms $GL_{n}(K)$ and the quadratic automorphism
(2) $(x_{1}+[x_{2},x_{3}],x_{2},\ldots,x_{n})$
if either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$, or $n\geq 5$ and $K$ is a field of arbitrary characteristic.
If $\phi=(f_{1},\ldots,f_{n}),\psi=(g_{1},\ldots,g_{n})\in\mathrm{Aut}(M_{n})$, then the composition $\phi\psi$ is defined as
$\phi\psi=(g_{1}(f_{1},\ldots,f_{n}),\ldots,g_{n}(f_{1},\ldots,f_{n})).$
We also adopt the standard notation $\phi,\psi]=\phi\psi\phi^{-1}\psi^{-1}$ for the commutator, and $\phi^{\psi}=\psi\phi\psi^{-1}$ for conjugation.
Let $0\neq z\in[M_{n},M_{n}]$. Then $\mathrm{ad}\,z:M_{n}\to M_{n}(m\mapsto[z,m])$ is a nilpotent derivation and $(\mathrm{ad}\,z)^{2}=0$. The exponential automorphism $\exp(\mathrm{ad}\,z)$ of $M_{n}$ can be written as
$\exp(\mathrm{ad}\,z)=\mathrm{id}+\mathrm{ad}\,z=(x_{1}+[z,x_{1}],\ldots,x_{n}+[z,x_{n}]).$
It is known that the exponential automorphism $\exp(\mathrm{ad}\,[x_{1},x_{2}])$ of $M_{3}$ is wild [14].
Recall that in 1993 Bryant and Drensky proved [7] that every automorphism of $M_{n}$ of rank $n\geq 4$ over a field $K$ of characteristic zero is approximately tame ( see the definitions in [47]); that is, it can be approximated by tame automorphisms with respect to power series topology [7, 47]. In contrast, all wild automorphisms of $M_{3}$ are absolutely wild [47], meaning that they cannot be approximated by tame automorphisms.
Free metabelian Lie algebras admit automorphisms of the form
$(x_{1},\ldots,x_{i-1},f_{i},x_{i+1},\ldots,x_{n}),\ 1\leq i\leq n,$
which include all elementary automorphisms but not necessarily elementary ones. They are called Chein automorphisms or one-row transformations [15, 42]. These automorphisms move only one variable and fix the others. The subgroup $\mathrm{ATAut}(M_{n})$ of $\mathrm{Aut}(M_{n})$ generated by all Chein automorphisms is called the almost tame automorphism group, and the elements of this subgroup is called almost tame automorphisms of $M_{n}$ [32, 34]. Thus
$\mathrm{TAut}(M_{n})\subseteq\mathrm{ATAut}(M_{n})\subseteq\mathrm{Aut}(M_{n}).$
Obviously, the group of almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated by all linear automorphisms $GL_{n}(K)$ and all Chein automorphisms of the form
(3) $(x_{1}+f,x_{2},\ldots,x_{n}),\ f\in[M_{n},M_{n}].$
Chein automorphisms of this form are described in Section 3.
The lowest-degree example of a non-elementary Chein automorphism is
(4) $(x_{1}+[[x_{2},x_{3}],x_{1}],x_{2},\ldots,x_{n}).$
This automorphism is wild if $n=3$ [32, 43], and is the first candidate for wildness for all $n\geq 4$, but it is still unknown whether it is tame or not. It was recently shown [36] that an analogue of this automorphism for free metabelian anticommutative algebras is absolutely wild for all $n\geq 3$.
In 2008, Roman’kov [43] proved that the group of automorphisms of $M_{3}$ cannot be finitely generated modulo the subgroup generated by all exponential and Chein automorphisms. Kabanov and Roman’kov [17] constructed strictly nontame primitive elements of $M_{3}$. Defining relations for Chein automorphisms of $M_{3}$ are given in [34].
We introduce the notion of the lower degree of an automorphism. Consider the standard degree grading of
$M_{n}=L_{1}\oplus L_{2}\oplus\ldots\oplus L_{k}\oplus\ldots,$
where $L_{k}$ is the linear span of all Lie monomials of degree $k$ in $x_{1},\ldots,x_{n}$. Every nonzero $m\in M_{n}$ can be uniquely written in the form
$m=l_{s}+l_{s+1}+\ldots+l_{t},$
where $l_{j}\in L_{j}$ for all $s\leq j\leq t$, $l_{s}\neq 0$, and $l_{t}\neq 0$. We call $\mathrm{ldeg}(m)=s$ the lower degree of $m$, while $\deg(m)=t$ denotes the degree of $m$. If $\mathrm{ldeg}(m)=\deg(m)=t$, then $m$ is homogeneous of degree $t$. Set also $\mathrm{ldeg}(0)=\infty$ and $\deg(0)=$-$\infty$.
Let
$\phi=(l_{1}+m_{1},\ldots,l_{n}+m_{n})$
be an automorphism of $M_{n}$ such that $l_{i}\in L_{1}$ and $m_{i}\in[M_{n},M_{n}]$ for all $i$. Set
$\mathrm{ldeg}(\phi)=\mathrm{min}\{\mathrm{ldeg}(m_{i})|1\leq i\leq n\}$
and
$\mathrm{deg}(\phi)=\mathrm{max}\{\mathrm{deg}(m_{i})|1\leq i\leq n\}.$
If $\mathrm{ldeg}(\phi)=\mathrm{deg}(\phi)=t$, then we say that $\phi$ is homogeneous of degree $t$. In particular, (2) is quadratic and (4) is cubic. Clearly, every nonlinear automorphism has a finite lower degree and degree.
We show that all Chein automorphisms of lower degree $\geq 4$ are tame. This means that all Chein automorphisms ”except” (4) are tame. We also show that all exponential automorphisms of lower degree $\geq 5$ are tame. These results contradict the results of [5, 39].
3. Chein automorphisms
In this section, using an analogue of the well known Magnus embedding [19, 23], we give another representation of the free metabelian algebra $M_{n}$, and give direct definition of the Fox derivatives for $M_{n}$. This construction for Lie algebras is a particular case of Shmel’kin’s wreath products [51].
Let $Y_{n}=M_{n}/[M_{n},M_{n}]$ be the abelian Lie algebra with a linear basis $y_{1},y_{2},\ldots,y_{n}$, where $y_{i}=x_{i}+[M_{n},M_{n}]$. The universal enveloping algebra $U(Y_{n})$ of $Y_{n}$ is the polynomial algebra $U=U(Y_{n})=K[y_{1},\ldots,y_{n}]$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Let
$U=U_{0}\oplus U_{1}\oplus\ldots\oplus U_{k}\oplus\ldots$
be the degree grading of $U$, i.e., $U_{k}$ is the linear span of all homogeneous monomials in $y_{1},y_{2},\ldots,y_{n}$ of degree $k$. If
$a=a_{s}+a_{s+1}+\ldots+a_{t}\in U,a_{j}\in U_{j},a_{s}\neq 0,a_{t}\neq 0,$
then $\deg(a)=t$ is the degree and $\mathrm{ldeg}(a)=s$ is the lower degree of $a$ in the variables $y_{1},y_{2},\ldots,y_{n}$. Set $\deg(0)=$-$\infty$ and $\mathrm{ldeg}(0)=\infty$ as usual.
Let $T_{n}$ be the free right $U$-module with basis $t_{1},\ldots,t_{n}$. Turn the direct sum
$M=Y_{n}\oplus T_{n}$
into a Lie algebra by
$[a+t,b+s]=tb-sa,$
where $a,b\in Y_{n}$ and $t,s\in T_{n}$. Then the subalgebra of $M$ generated by
$x_{i}=y_{i}+t_{i},\ \ 1\leq i\leq n,$
is the free metabelian Lie algebra with free generators $x_{1},\ldots,x_{n}$ [19, 23, 51]. We identify this algebra with $M_{n}$.
Notice that $M_{n}^{\prime}=[M_{n},M_{n}]\subseteq T_{n}$ and for any $m\in M_{n}^{\prime}$ and $y_{i}$ we get
$my_{i}=[m,x_{i}].$
Every element $m\in M_{n}^{\prime}$ can be written in the form
$m=\sum_{1\leq i<j\leq n}[x_{i},x_{j}]a_{ij},\ \ a_{ij}\in U.$
Let $f\in M_{n}$ be an arbitrary element $M_{n}$. Then $f$ is uniquely represented as
$f=y+t,y\in Y_{n},t\in T_{n}.$
Moreover, $t$ is uniquely represented as
$t=t_{1}d_{1}+\ldots+t_{n}d_{n},\ \ \ \ d_{1},\ldots,d_{n}\in K[y_{1},\ldots,y_{n}]=U.$
Set $d_{i}=\frac{\partial f}{\partial x_{i}}$ for all $1\leq i\leq n$ and
$\partial(f)=(d_{1},\ldots,d_{n})^{t}=\big(\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}\big)^{t},$
where $t$ means the transpose. We call $\frac{\partial f}{\partial x_{1}},\ldots,\frac{\partial f}{\partial x_{n}}$ the Fox derivatives of $f$ (see [53, 56]).
Let $e_{1},e_{2},\ldots,e_{n}$ be the columns of the identity matrix $I=I_{n}$ of order $n$.
The statement of the next lemma is well known [56, 58].
Lemma 1.
Let $a=(a_{1},a_{2},\ldots,a_{n})^{t}\in(U^{n})^{t}$ be an arbitrary column of dimension $n$ over $U$. Then the following conditions are equivalent.
$(a)$ $a=\partial(f)$ for some $f\in[M_{n},M_{n}]$;
$(b)$ $Ya=y_{1}a_{1}+\ldots+y_{n}a_{n}=0$ where $Y=(y_{1},\ldots,y_{n})$;
$(c)$ $a=\sum_{i<j}(e_{i}y_{j}-e_{j}y_{i})a_{ij},\ a_{ij}\in U$.
Proof. If $i<j$ and $a\in U$, then
$\partial([x_{i},x_{j}]a)=e_{i}y_{j}a-e_{j}y_{i}a=(e_{i}y_{j}-e_{j}y_{i})a.$
Consequently, $(a)$ implies $(c)$. Obviously, $(c)$ implies $(b)$ since $Y(e_{i}y_{j}-e_{j}y_{i})=0$ for all $i<j$.
Suppose that $(b)$ holds. Let $a_{i}=-y_{n}b_{i}+a_{i}^{0}$, where $a_{i}^{0}\in K[y_{1},\ldots,y_{n-1}]$ for all $1\leq i\leq n-1$. Then $(b)$ implies that $a_{n}=y_{1}b_{1}+\ldots+y_{n-1}b_{n-1}$. Set $g=[x_{n},x_{1}]b_{1}+\ldots+[x_{n},x_{n-1}]b_{n-1}$. We get $\frac{\partial g}{\partial x_{n}}=a_{n}$ and the last component of $a^{\prime}=a-\partial(g)$ is zero. Moreover, if $a^{\prime}=(a_{1}^{\prime},\ldots,a_{n-1}^{\prime},0)$ then $y_{1}a_{1}^{\prime}+\ldots+y_{n-1}a_{n-1}^{\prime}=0$. Leading an induction on the number of nonzero components of $a$, we may assume that there exists $h\in[M_{n},M_{n}]$ such that $a^{\prime}=\partial(h)$. Then $a=\partial(g+h)$. $\Box$
Every endomorphism $\phi$ of $M_{n}$ induces an endomorphism of $Y_{n}=M_{n}/[M_{n},M_{n}]$, which can be uniquely extended to an endomorphism of $U=U(Y_{n})$. This endomorphism will be denoted again by $\phi$ for simplicity of notation. If $f\in U$, then the image $\phi(f)$ will be denoted by $f^{\phi}$ for convenience.
We define the Jacobian matrix of $\phi=(f_{1},\ldots,f_{n})$ by
$J(\phi)=\left[\frac{\partial f_{j}}{\partial x_{i}}\right]_{1\leq i,j\leq n}=\left[\partial(f_{1})\ \ldots\ \partial(f_{n})\right].$
If $\phi=(f_{1},\ldots,f_{n})$ and $\psi=(g_{1},\ldots,g_{n})$ then
$\phi\psi(x_{i})=g_{i}(f_{1},\ldots,f_{n})$
for all $1\leq i\leq n$. The corresponding Chain Rule gives [53, 56, 58] that
$J(\phi\psi)=J(\phi)J(\psi)^{\phi}.$
Consequently, if $\phi$ is an automorphism, then $J(\phi)$ is invertible. The converse is also true [53, 56].
Lemma 2.
An endomorphism $\phi$ of $M_{n}$ is an automorphism if and only if the Jacobian matrix $J(\phi)$ is invertible.
Proof. Suppose that $J(\phi)$ is invertible. Without loss of generality, we may assume that
$\phi=(x_{1}+f_{1},\ldots,x_{n}+f_{n}),\ \ f_{1},\ldots,f_{n}\in[M_{n},M_{n}].$
Then
$J(\phi)=I+A,\ \ A=[\partial(f_{1}),\ldots,\partial(f_{n})].$
We get
$J(\phi)^{-1}=I+B,B=\sum_{i\geq 1}(-A)^{i}.$
By Lemma 1, we have $YA=0$. Consequently, $YB=0$. Again, by Lemma 1, there exist $g_{1},\ldots,g_{n}\in[M_{n},M_{n}]$ such that
$B=[\partial(g_{1}),\ldots,\partial(g_{n})].$
Let $\psi=(x_{1}+g_{1},\ldots,x_{n}+g_{n})$. Then
$J(\psi)=I+B=J(\phi)^{-1}$
and
$J(\phi\psi)=J(\phi)J(\psi)^{\phi}=J(\phi)J(\psi)=I.$
Consequently, $\phi\psi=\mathrm{Id}$ and $\phi$ is an automorphism. $\Box$
We now give a description of Chein automorphisms of the form (3).
Lemma 3.
An endomorphism of $M_{n}$ of the form (3) is an automorphism if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
Proof. Let $\phi=(x_{1}+f,x_{2},\ldots,x_{n})$ where $f\in[M_{n},M_{n}]$. The Jacobian matrix of $\phi$ is
$J(\phi)=I+[\partial(f)\ 0\ldots 0].$
This matrix is invertible if and only if $1+\frac{\partial f}{\partial x_{1}}$ is invertible in $U$ or, equivalently, $\frac{\partial f}{\partial x_{1}}=0$. We want to show that $\frac{\partial f}{\partial x_{1}}=0$ if and only if $f$ belongs to the ideal generated by all $x_{s},x_{t}]$, where $s,t\neq 1$.
The basis (1) of $M_{n},M_{n}]$ is convenient to work with $\frac{\partial f}{\partial x_{n}}=0$. Any $f\in[M_{n},M_{n}]$ can be written in the form
$f=f^{\prime}+[x_{n},x_{1}]a_{1}+[x_{n},x_{2}]a_{2}+\ldots+[x_{n},x_{n-1}]a_{n-1},$
where $f^{\prime}$ belongs to the ideal generated by all $x_{s},x_{t}]$ with $s,t\neq n$ and $a_{i}\in K[y_{i},\ldots,y_{n}]$ for all $i$. Since $\frac{\partial f^{\prime}}{\partial x_{n}}=0$ it follows that
$\frac{\partial f}{\partial x_{n}}=y_{1}a_{1}+y_{2}a_{2}+\ldots+y_{n-1}a_{n-1}.$
If
$y_{1}a_{1}+y_{2}a_{2}+\ldots+y_{n-1}a_{n-1}=0$
then the substitution $y_{1}=\ldots=y_{n-2}=0$ gives that $a_{n-1}=0$. Repeated substitution $y_{1}=\ldots=y_{n-3}=0$ gives $a_{n-2}=0$. Continuing the same discussions, we get $a_{n-1}=a_{n-2}=\ldots=a_{1}=0$. Therefore $\frac{\partial f}{\partial x_{n}}=0$ if and only if $f=f^{\prime}$. $\Box$
4. Tame automorphisms
In this section we always assume that either $n\geq 4$ and $K$ is a field of characteristic $\neq 3$ or $n\geq 5$.
Theorem 1.
Every Chein automorphism of $M_{n}$ of lower degree $\geq 4$ is tame.
Proof. It suffices to prove that every Chein automorphism of the form
$C(a)=(x_{1}+[x_{2},x_{3}]a,x_{2},\ldots,x_{n}),\ a\in U,$
with $\mathrm{ldeg}(a)\geq 2$ is tame. Since
$C(a+b)=C(a)C(b),\ \ a,b\in U,$
we may assume that $a=\gamma y_{1}^{i_{1}}y_{2}^{i_{2}}\ldots y_{n}^{i_{n}}$, $0\neq\gamma\in K$. If $i_{1}=0$ then $C(a)$ is elementary. Assume that $i_{1}\geq 1$.
Case 1. Suppose that $i_{j}\geq 1$ for some $2\leq j\leq n-1$. Consider the elementary automorphisms
$\phi=(x_{1},\ldots,x_{n-1},x_{n}+[x_{2},x_{3}]y_{1}^{i_{1}}\ldots y_{j}^{i_{j}-1}),$
$\psi=(x_{1}-\gamma[x_{j},x_{n}]y_{j+1}^{i_{j+1}}\ldots y_{n}^{i_{n}},x_{2},\ldots,x_{n}).$
Direct calculations give that $\phi,\psi]=C(a)$ is tame.
Case 2. Let $i_{2}=i_{3}=\ldots=i_{n-1}=0$ and $i_{n}\neq 0$. Then $a=\gamma y_{1}^{i_{1}}y_{n}^{i_{n}}$. We can make $\gamma=1$ by conjugating $C(a)$ with $(x_{1},\gamma^{-1}x_{2},x_{3}\ldots,x_{n})$. Then assume that $a=y_{1}^{i_{1}}y_{n}^{i_{n}}$. We have
(5) $[x_{2},x_{3}]a=[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1}$
by the Jacobi identity. If $i_{n}\geq 2$ then
$\phi=(x_{1}+[x_{2},x_{n}]y_{1}^{i_{1}}y_{3}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{3}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(3n)}$
and
$\psi=(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{2}y_{n}^{i_{n}-1},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2}^{i_{n}-1}y_{n},x_{2},\ldots,x_{n})^{(2n)}$
are tame by Case 1. Therefore $C(a)=\phi\psi$ is tame.
If $i_{n}=1$, then set $\alpha=(x_{1},x_{2}+x_{n},x_{3},\ldots,x_{n})$. We have
$\tau=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})$
is tame by Case 1. Direct calculations give that
$[\alpha,\tau]=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})$
is tame. Since the second factor
$(x_{1}+[x_{n},x_{3}]y_{1}^{i_{1}}y_{n},x_{2},\ldots,x_{n})=(x_{1}+[x_{2},x_{3}]y_{1}^{i_{1}}y_{2},x_{2},\ldots,x_{n})^{(2n)}=\tau^{(2n)}$
of this product is also tame it follows that its first factor
$\sigma=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
is tame. By (5), we get
$3[x_{2},x_{3}]a=([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}}.$
Consequently,
$C(3a)=(x_{1}+([x_{2},x_{3}]y_{n}+[x_{n},x_{3}]y_{2})y_{1}^{i_{1}},x_{2},\ldots,x_{n})$
$\times(x_{1}+([x_{2},x_{3}]y_{n}+[x_{2},x_{n}]y_{3})y_{1}^{i_{1}},x_{2},\ldots,x_{n})=\sigma(\sigma^{-1})^{(23)}$
is tame. If the characteristic of $K$ is not $3$, then its conjugation by $(x_{1},3x_{2},x_{3},\ldots,x_{n})$ gives $C(a)$.
If $n\geq 5$, then conjugating $C(a)$ by $(4n)$, we can make $i_{4}>0$. Case 1 gives that $C(a)$ is tame again.
Case 3. Let $i_{2}=i_{3}=\ldots=i_{n-1}=i_{n}=0$. In this case $a=y_{1}^{i_{1}}$ and $s=i_{1}\geq 2$ since $\mathrm{ldeg}(a)\geq 2$. Case 2 gives that
$\phi=(x_{1}+[x_{2},x_{3}]y_{1}^{s-1}y_{n},x_{2},\ldots,x_{n})$
is tame. Set $\beta=(x_{1},,\ldots,x_{n-1},x_{n}+x_{1})$. Direct calculations give that
$[\beta,\phi]=C(a)(x_{1},\ldots,x_{n-1},x_{n}-[x_{2},x_{3}]y_{1}^{s-1}(y_{1}+y_{n})$
$=C(a)(x_{1}-[x_{2},x_{3}]y_{n}^{s},\ldots,x_{n})^{(1n)}(x_{1}-[x_{2},x_{3}]y_{1}y_{n}^{s-1},x_{2},\ldots,x_{n})^{(1n)}.$
Since the second factor of this product is tame and the third factor is tame by Case 2 it follows that $C(a)$ is tame. $\Box$
This theorem contradicts the results of [39].
Corollary 1.
The group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4).
Proof. The group of all tame automorphisms $\mathrm{TAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) [35]. Obviuosly, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by Chein automorphisms $C(a)$. We have
$C(a)=C(a_{0})C(a_{1})\ldots C(a_{k}),$
where $a_{i}\in U_{i}$ for all $0\leq i\leq k$. We have $C(a_{0})$ is elementary since $a_{0}\in K$ and $C(a_{2}),\ldots,C(a_{k})$ are tame by Theorem 1.
Let
$a_{1}=\alpha_{1}y_{1}+\alpha_{2}y_{2}+\ldots+\alpha_{n}y_{n}.$
Then
$C(a_{1})=C(\alpha_{1}y_{1})C(\alpha_{2}y_{2})\ldots C(\alpha_{n}y_{n}).$
All $C(\alpha_{2}y_{2}),\ldots,C(\alpha_{n}y_{n})$ are again tame.
Consider $C(\alpha_{1}y_{1})$. Conjugating this by $(x_{1},\alpha_{1}^{-1}x_{2},x_{3},\ldots,x_{n})$ we can make $\alpha_{1}=1$ if $\alpha_{1}\neq 0$.
Consequently, the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ modulo $\mathrm{TAut}(M_{n})$ is generated by $C(y_{1})$. Note that $C(y_{1})$ is the cubic automorphism (4). Therefore the group of all almost tame automorphisms $\mathrm{ATAut}(M_{n})$ is generated, modulo all linear automorphisms, by the quadratic automorphism (2) and the cubic automorphism (4). $\Box$
Lemma 4.
Every automorphism of $M_{n}$ of the form
$D(a)=(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
is tame if $\mathrm{ldeg}(a)\geq 2$.
Proof. By Theorem 1, if $\mathrm{ldeg}(b)\geq 1$, then
$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
is tame. Let
$\psi=(x_{1},x_{2},x_{3}+[x_{1},x_{2}],x_{4},\ldots,x_{n}).$
Then
$[\phi,\psi]=(x_{1}+[x_{1},x_{2}]y_{1}y_{4}b,x_{2}+[x_{1},x_{2}]y_{2}y_{4}b,x_{3},\ldots,x_{n})=D(y_{4}b)$
is tame.
Consequently, the statement of the lemma is true if $y_{4}$ divides $a$. Obviously, the same is true if $y_{i}$ divides $a$ for all $3\leq i\leq n$. Since
$D(a_{1}+a_{2})=D(a_{1})+D(a_{2}),\ \ a_{1},a_{2}\in U,$
we may assume that $a\in K[y_{1},y_{2}]$. Let $\alpha=(x_{1},x_{2},x_{3},x_{4}+x_{1},x_{5},\ldots,x_{n})$. If $b\in K[y_{1},y_{2}]$ then
$D(y_{4}b)^{\alpha}=(x_{1}+[x_{1},x_{2}]y_{1}(y_{4}+y_{1})b,x_{2}+[x_{1},x_{2}]y_{2}(y_{4}+y_{1})b,$
$x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
$=D(y_{4}b)D(y_{1}b)(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n}).$
Since $D(y_{4}b)$ is tame and
$(x_{1},x_{2},x_{3},x_{4}-[x_{1},x_{2}]y_{2}(y_{4}+y_{1}),x_{5},\ldots,x_{n})$
is tame by Theorem 1, this implies that $D(y_{1}b)$ is also tame. Therefore the statement of the lemma is true if $a\in K[y_{1},y_{2}]$ and $y_{1}$ divides $a$.
Changing $\alpha$ to $\beta=(x_{1},x_{2},x_{3},x_{4}+x_{2},x_{5},\ldots,x_{n})$ and repeating the same discussions as above, we can show that $D(y_{2}b)$ is also tame. $\Box$
Theorem 2.
Every exponential automorphism of $M_{n}$ of lower degree $\geq 5$ is tame.
Proof. Let
$E(m)=\exp(\mathrm{ad}(m))=(x_{1}+[m,x_{1}],x_{2}+[m,x_{2}],\ldots,x_{n}+[m,x_{n}])$
be an exponential automorphism with $\mathrm{ldeg}(m)\geq 4$. Since
$E(m_{1}+m_{2})=E(m_{1})+E(m_{2}),$
we may assume that $m=[x_{1},x_{2}]a$ with $\mathrm{ldeg}(a)\geq 2$. Then
$E(m)=E([x_{1},x_{2}]a)=D(a)\phi_{3}\ldots\phi_{n},$
where
$\phi_{i}=(x_{1},\ldots,x_{i-1},x_{i}+[x_{1},x_{2}]y_{i}a,x_{i+1},\ldots,x_{n}),\ 3\leq i\leq n.$
All $\phi_{i}$ are tame by Theorem 1 and $D(a)$ is tame by Lemma 4. Consequently, $E(m)$ is tame. $\Box$
This theorem contradicts the main result of [5].
5. Almost tame automorphisms
In this section we prove that a huge class of automorphisms of $M_{n}$ that moves only two variables $x_{1}$ and $x_{2}$ are almost tame. We always assume that $n\geq 4$ and $K$ is an arbitrary field. Since all Chein automorphisms are automatically included into $\mathrm{ATAut}(M_{n})$ by the definition, we don’t need any restrictions to characteristic of the field from the preceeding section.
First we give minor extensions of Lemma 4 and Theorem 2 for almost tame automorphisms.
Lemma 5.
Every automorphism of the form
$D(a)=(x_{1}+[x_{1},x_{2}]y_{1}a,x_{2}+[x_{1},x_{2}]y_{2}a,x_{3},\ldots,x_{n})$
is tame if $\mathrm{ldeg}(a)\geq 1$.
Proof. For any $b\in U$ we have
$\phi=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
$=(x_{1}-[x_{3},x_{4}]y_{1}b,x_{2}-[x_{3},x_{4}]y_{2}b,x_{3},\ldots,x_{n})$
is almost tame. Repeating the proof of Lemma 4 with this data, we get that $D(a)$ is almost tame for all $a$ with $\mathrm{ldeg}(a)\geq 1$. $\Box$
Corollary 2.
Every exponential automorphism of lower degree $\geq 4$ is almost tame.
Proof. Consider the decomposition of the exponential automorphism $E([x_{1},x_{2}]a)$ from the proof of Theorem 2. Note that $\mathrm{ldeg}(E([x_{1},x_{2}]a))\geq 4$ implies $\mathrm{ldeg}(a)\geq 1$. Consequently, $D(a)$ is almost tame by Lemma 5. Since all $\phi_{i}$ are Chein automorphisms it follows that $E([x_{1},x_{2}]a)$ is almost tame. Consequently, every exponential automorphism of lower degree $\geq 4$ belongs to $\mathrm{ATAut}(M_{n})$. $\Box$
Lemma 6.
Every automorphism of the form
$A(h,g)=(x_{1}+[x_{1},x_{n}]hg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]h-[x_{2},x_{n}]hg,x_{3},\ldots,x_{n}),$
where $g,h\in U$, is almost tame.
Proof. $(a)$ Let $\alpha=(x_{1}-\lambda x_{2},x_{2},\ldots,x_{n})$. Then
$A(h,g)^{\alpha}=A(h^{\alpha},g^{\alpha}-\lambda).$
This means that $A(h,g)$ is almost tame if and only if $A(h^{\alpha},g^{\alpha}-\lambda)$ is almost tame. Using this we may assume that $g(0,\ldots,0)=0$.
$(b)$ Consider the almost tame automorphisms
$\phi=(x_{1}-[x_{2},x_{3}]a,x_{2},\ldots,x_{n})$
and
$\psi=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2},\ldots,x_{n})(x_{1},x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
$=(x_{1}+[x_{3},x_{n}]y_{2}ah(g-y_{3}a),x_{2}-[x_{3},x_{n}]y_{2}ah,x_{3},\ldots,x_{n})$
Direct calculations give that
$A(h,g)^{\phi}=A(h,g-y_{3}a)\psi.$
Consequently, $A(h,g)$ is almost tame if and only if $A(h,g-y_{3}f)$ is almost tame. Of course, the same trick can be done with any variable $y_{i}$ if $3\leq i\leq n-1$. Therefore, replacing $g$ by $g-y_{i}f$ for all $3\leq i\leq n-1$, we may assume that
$g=G(y_{1},y_{2})-y_{n}b,\ \ b\in U.$
$(c)$ Now we show how to eliminate $y_{n}b$. First we replace $g$ to
$g+y_{3}b=G(y_{1},y_{2})+y_{3}b-y_{n}b$
using $(b)$. If $\beta=(x_{1},x_{2},x_{3}+x_{n},x_{4},\ldots,x_{n})$, then it is easy to check that
$A(h,g+y_{3}b)^{\beta}=A(h^{\beta},(g+y_{3}b)^{\beta}).$
We have
$(g+y_{3}b)^{\beta}=g^{\beta}+(y_{3}+y_{n})b^{\beta}$
$=G(y_{1},y_{2})-y_{n}b^{\beta}+(y_{3}+y_{n})b^{\beta}=G(y_{1},y_{2})+y_{3}b^{\beta}.$
Using $(b)$ again, we may assume that
$g=G(y_{1},y_{2}).$
(d) Let $c$ be a polynomial in $y_{1},y_{2}$. Then $g$ can be replaced by $g+y_{i}c$, where $i=1,2$. For this, we first replace $g$ by $g+y_{3}c$ using $(b)$, i.e., we can assume that $g=G(y_{1},y_{2})+y_{3}c$. If $\gamma=(x_{1},x_{2},x_{3}+x_{1},x_{4},\ldots,x_{n})$, then
$A(h,g)^{\gamma}=\sigma A(h^{\gamma},g^{\gamma}),$
where
$\sigma=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
$=(x_{1},x_{2},x_{3}-[x_{1},x_{n}]h^{\gamma}g^{\gamma},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}-[x_{2},x_{n}]h^{\gamma}(g^{\gamma})^{2},x_{4},\ldots,x_{n})$
is an almost tame automorphism. Consequently, $A(h,g)$ is almost tame if and only if $A(h^{\gamma},g^{\gamma})$ is almost tame. Note that
$g^{\gamma}=G(y_{1},y_{2})+y_{3}c+y_{1}c.$
Using $(b)$, again we can replace $g^{\gamma}$ by
$G(y_{1},y_{2})+y_{1}c.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{1}c$ for any $c\in K[y_{1},y_{2}]$.
If $\delta=(x_{1},x_{2},x_{3}+x_{2},x_{4},\ldots,x_{n})$ and $g=G(y_{1},y_{2})+y_{3}c$, then
$A(h,g)^{\delta}=\tau A(h^{\delta},g^{\delta}),$
where
$\tau=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
$=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]h^{\delta},x_{4},\ldots,x_{n})(x_{1},x_{2},x_{3}+[x_{2},x_{n}]h^{\delta}g^{\delta},x_{4},\ldots,x_{n})$
is again almost tame automorphism. In this case we have
$g^{\delta}=G(y_{1},y_{2})+y_{3}c+y_{2}c.$
Using $(b)$, we can replace $g^{\delta}$ by
$G(y_{1},y_{2})+y_{2}c.$
This means that we can replace $g=G(y_{1},y_{2})$ by $G(y_{1},y_{2})+y_{2}c$ for any $c\in K[y_{1},y_{2}]$.
Thus we can make $g=0$. Observe that
$A(h,0)=(x_{1},x_{2}-[x_{1},x_{n}]h,x_{3},\ldots,x_{n})$
is a Chein automorphism. $\Box$
Theorem 3.
Every automorphism of the form
$B(h,f,g)=(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{2},x_{n}]hfg,x_{3},\ldots,x_{n})$
is almost tame.
Proof. Consider the Chein automorphisms
$\phi_{1}=(x_{1},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{3},x_{n}]hf^{2}g,x_{3},\ldots,x_{n}),$
$\psi_{1}=(x_{1}+[x_{2},x_{n}]hg^{2}-[x_{3},x_{n}]hfg^{2},x_{2},\ldots,x_{n})$
and the automorphisms
$\phi_{2}=(x_{1}+[x_{1},x_{n}]hfg+[x_{3},x_{n}]hfg^{2},x_{2},x_{3}-[x_{1},x_{n}]hf-[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n}),$
$\psi_{2}=(x_{1},x_{2}-[x_{2},x_{n}]hfg+[x_{3},x_{n}]hf^{2}g,x_{3}-[x_{2},x_{n}]hg+[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n}).$
Note that
$(\phi_{2})^{(23)}=A((hf)^{(23)},g^{(23)})\ \hbox{and}\ (\psi)^{132)}=A((hg)^{(132)},-f^{(132)})$
are almost tame by Lemma 6. Then $\phi_{2}$ and $\psi_{2}$ are also almost tame. Conseqently,
$\phi=\phi_{1}\phi_{2}=(x_{1}+[x_{1},x_{n}]hfg+[x_{3},x_{n}]hfg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{3},x_{n}]hf^{2}g,$
$x_{3}-[x_{1},x_{n}]hf-[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n})$
and
$\psi=(x_{1}+[x_{2},x_{n}]hg^{2}-[x_{3},x_{n}]hfg^{2},x_{2}-[x_{2},x_{n}]hfg+[x_{3},x_{n}]hf^{2}g,$
$x_{3}-[x_{2},x_{n}]hg+[x_{3},x_{n}]hfg,x_{4},\ldots,x_{n})$
are almost tame. Direct calculations give that
$\phi\psi=(x_{1}+[x_{1},x_{n}]hfg+[x_{2},x_{n}]hg^{2},x_{2}-[x_{1},x_{n}]hf^{2}-[x_{2},x_{n}]hfg,$
$x_{3}-[x_{1},x_{n}]hf-[x_{2},x_{n}]hg,x_{4},\ldots,x_{n}).$
Then
$B(h,f,g)=\tau\phi\psi,$
where
$\tau=(x_{1},x_{2},x_{3}+[x_{1},x_{n}]hf+[x_{2},x_{n}]hg,x_{4},\ldots,x_{n}).$
Note that $\tau$ is a Chein automorphism. Hence $B(h,f,g)$ is almost tame. $\Box$
Acknowledgments
The author would like to thank the Max Planck Institute für Mathematik for its hospitality and excellent working conditions, where some part of this work has been done. The author thanks Professors V. Dotsenko, M. Jibladze, L. Makar-Limanov, and I. Shestakov for useful discussions.
The project is supported by the grant AP14872073 of the Ministry of Education and Science of the Republic of Kazakhstan.